# Your Token Becomes Worthless: Unveiling Rug Pull Schemes in Crypto Token via Code-and-Transaction Fusion Analysis

### v1

@@ Abstract.
In this paper, we propose RPHunter, a novel technique that integrates code and transaction for Rug Pull detection. First, RPHunter establishes declarative rules and performs flow analysis to extract code risk information, further constructing a semantic risk code graph (SRCG). Meanwhile, to leverage transaction information, RPHunter formulates dynamic token transaction activities as a token flow behavior graph (TFBG) in which nodes and edges are characterized from network structure and market manipulation perspectives. Finally, RPHunter employs graph neural networks to extract complementary features from SRCG and TFBG, integrating them through an attention fusion model to enhance the detection of Rug Pull. We manually analyzed 645 Rug Pull incidents from code and transaction aspects and constructed a ground-truth dataset. We evaluated RPHunter on our dataset, achieving a precision of 95.3%, a recall of 93.8% and an F1 score of 94.5%, which highlights superior performance compared to existing state-of-the-art methods. Furthermore, when applied to the real-world scenarios, RPHunter has identified 4801 Rug Pull tokens, achieving a precision of 91%.
copyright: nonejournalyear: 2025publicationmonth: March
1. Introduction
@@
Blockchain technology has revolutionized the financial landscape, enabling direct peer-to-peer transactions and financial services without centralized intermediaries. A diverse range of decentralized applications, including decentralized exchanges, lending platforms, and prediction markets, have been built on it. At the time of writing, the global crypto market cap has reached \$3.07 trillion (Coi, 2025). However, the rapid expansion and the inherent openness of DeFi also pose considerable security challenges. Among these, Rug Pulls have emerged as a prominent form of scam, leading to significant financial losses, with approximately \$85.4 million in 2024 alone (Certik, 2025).
@@
Rug Pulls are a pernicious form of scam prevalent within the DeFi ecosystem, where scammers lure investors with seemingly promising tokens. Once a substantial amount of funds is collected, these scammers abruptly withdraw the funds and abandon the project entirely, leaving investors with worthless tokens. To detect such scams, researchers have proposed various methods (Ma et al., 2023; Lin et al., 2024; Zhou et al., 2024; Xia et al., 2021; Mazorra et al., 2022), broadly classified into rule-based methods (Lin et al., 2024; Zhou et al., 2024) and learning-based methods (Xia et al., 2021; Mazorra et al., 2022). Rule-based methods typically use static analysis to identify vulnerabilities in token contracts, requiring domain knowledge of known Rug Pull patterns. Despite their promise in detecting these known patterns, these methods are prone to high false negatives due to their limited detection scope. In contrast, learning-based methods employ statistical features from token transactions and opcode data to train machine learning models for detection. However, these methods struggle to identify coordinated manipulation involving multiple addresses and detect Rug Pulls before they occur, suffering from high false positives and poor practicality.
@@
The continuous evolution of real-world Rug Pulls makes them increasingly complex and diverse, posing significant challenges to existing methods. (1) Real-world Rug Pull schemes are constantly evolving and it is a complex interplay between malicious code and suspicious behaviors, yet there is still a lack of research that investigates the underlying mechanisms. Although some studies (Lin et al., 2024; Zhou et al., 2024) have partially uncovered the mechanisms, a comprehensive empirical study focusing on the mechanisms of real-world Rug Pull cases is still missing. (2) In real-world scenarios, Rug Pull schemes typically begin with the injection of malicious code and then facilitate market manipulation through suspicious transaction behaviors. Relying solely on code or transaction information to detect Rug Pulls can lead to false negatives and false positives. However, existing methods fail to effectively organize and leverage information from both code and transaction perspectives.
To address these challenges, we have conducted an empirical study to uncover the underlying mechanisms of Rug Pull. We collected 645 real-world Rug Pull incidents from security platforms and manually analyzed them from code and transaction aspects. Compared to the most recent studies (Zhou et al., 2024; Lin et al., 2024) with datasets of 103 and 201 samples respectively, we have constructed, to the best of our knowledge, the largest manually analyzed Rug Pull dataset and summarized the Rug Pull procedure and Rug Pull risk features.
@@
We evaluated RPHunter on our ground-truth dataset. We compared the performance of RPHunter with five state-of-the-art methods. The results demonstrate that RPHunter detects Rug Pull tokens with a precision of 95.3%, a recall of 93.8%, and an F1 score of 94.5%, significantly outperforming existing methods. Additionally, to assess its real-world applicability, we applied RPHunter to newly created Ethereum Mainnet crypto tokens. RPHunter identified 4801 Rug Pull tokens. The precision of real-world evaluation reaches 91%. A prototype of RPHunter, all of our datasets, and experimental results are available online (RPhunter, 2025).
@@
  • We construct a manually analyzed Rug Pull dataset, comprising 645 Rug Pulls, and summarize the typical Rug Pull procedure along with risk features from code and transaction perspectives.
@@
  • To overcome the limitations of existing methods that fail to fully utilize contract code and transaction information, we propose RPHunter, a novel method that integrates the complementary information from code risk and transaction behavior, utilizing GNNs for graph feature learning to detect Rug Pull schemes.
@@
  • Experimental results show that RPHunter outperforms the current SOTA methods on the ground-truth dataset with a precision of 95.3%, a recall of 93.8%, and an F1 score of 94.5%. Additionally, we have identified 4801 Rug Pull tokens in real-world crypto tokens, achieving a precision of 91%.
2. Background
2.1. Smart Contract and Transaction
@@
In the blockchain, smart contracts are self-executing programs with predefined rules and logic encoded in their code. Smart contracts are typically written in a Turing-complete programming language (e.g., Solidity, Viper). To deploy smart contracts on the blockchain, they need to be compiled into bytecode and then submitted to the blockchain with transactions. Once deployed, smart contracts are immutable and can be accessed through public functions, serving as entry points for interactions with external owned accounts (EOAs) or other contracts. Transactions are essential as the medium for facilitating interactions within the blockchain network, acting as the vehicle for operations such as contract creation and invocation.
2.2. Token and Token Standard
@@
Tokens are a kind of cryptocurrency implemented by smart contracts which manage internal book-keeping to record the information of token holders (e.g., their addresses and shares), and support token-related activities such as token balance inquiry and token transfer. Token standards specify a unified protocol for token creation, distribution, and interaction, ensuring compatibility and functionality across various platforms. A token smart contract should implement standard interfaces and emit standard events defined by token standards, otherwise, the token cannot be purchased or sold on the blockchain. Tokens are mainly classified into fungible tokens and non-fungible tokens, which follow the ERC-20 and ERC-721 standards or derived standards. For example, the ERC-20 standard defines 6 standard interfaces and 2 standard events for transferring tokens, retrieving account balances, and approving the use of tokens, alongside events for logging transactions.
2.3. Rug Pull
@@
Rug Pulls are a common form of scam within the DeFi ecosystem, often occurring in Decentralized Exchanges (DEXs). Unlike centralized exchanges, which rely on matching buyers and sellers, DEXs use liquidity pools for trading. Liquidity pools are a fundamental mechanism in DEXs where tokens are paired with reputable assets (e.g., WETH), allowing users to trade tokens directly from the pool without intermediaries. The liquidity pool dynamically adjusts token prices based on the principles of supply and demand. A typical Rug Pull begins with scammers deploying a scam token and creating a liquidity pool using the token paired with a reputable asset. As users purchase the token, its price will rise, creating an illusion of legitimacy and attracting more investors. Once the token price increases significantly, the scammers exploit their control over the liquidity pool or token contract by swiftly exchanging a large number of scam tokens for valuable tokens. This sudden exchange drains the liquidity, leaving investors with worthless tokens.
3. Real-world Rug Pull Study
3.1. Data Collection and Analysis
3.1.1. Rug Pull Incidents Collection
@@
To conduct an empirical study on Rug Pull schemes, we extensively gathered as many Rug Pull incidents as possible from blockchain security platforms including De.Fi (De., 2025), PeckShield (Pec, 2025), Certik (Cer, 2025), SlowMist (Slo, 2025), Beosin (Beo, 2025), BlockSec (blo, 2025), RugDoc (Rug, 2025), Chainalysis (Cha, 2025) and Solidus Labs (Sol, 2025). Ultimately, we collected a total of 1048 Rug Pull incidents that occurred before May 2024, encompassing multiple blockchain networks, including ETH (Eth, 2025), BSC (BSC, 2025), FTM (Fan, 2025), BASE (BAS, 2025), Arbitrum (Arb, 2025) and Polygon (Pol, 2025).
@@
After analyzing these Rug Pull incidents, we excluded 131 tokens without source code, hindering our precise analysis of these cases, and 272 tokens without transaction history, which suggests that these projects did not result in actual Rug Pulls. Finally, we obtained 645 Rug Pull incidents. Compared to the most recent studies (Zhou et al., 2024; Lin et al., 2024) with datasets of 103 and 201 samples, we constructed, to the best of our knowledge, the largest manually analyzed Rug Pulls dataset (RPhunter, 2025). While some studies (Xia et al., 2021; Mazorra et al., 2022) have constructed datasets of thousands of Rug Pull cases, they have two limitations: (1) these cases lack manual analysis and verification (e.g., code risk analysis), (2) rely on heuristic rules to select these cases, making them susceptible to false positives.
3.1.2. Incidents Analysis
@@
To better understand Rug Pull schemes, we undertook a comprehensive manual analysis of the incidents. For each Rug Pull incident, we conducted a detailed examination of the tokens, focusing on both the source code and the transaction history to uncover the underlying mechanisms of each Rug Pull. To ensure a high level of accuracy, we recruited three experts with two years of experience in smart contract security research and conducted a two-round analysis of these Rug Pull incidents.
In the first round, we independently reviewed token contracts and transaction records, identifying potential code risks and unusual transaction behaviors. Each expert documented the root causes of the incidents, followed by a consensus-based voting session to determine the most plausible cause for each Rug Pull. The second round synthesized the insights from the initial analysis. We collaboratively reviewed the identified root causes, formulating a generalized Rug Pull procedure (Section 3.2), which captures the essence of these fraudulent activities. We further summarized the Rug Pull risk features that emerge from code risk (Section 3.3.1) and transaction behavior (Section 3.3.2) two perspectives.
3.2. Rug Pull Procedure
After a systematic analysis of the Rug Pull procedure based on these real-world incidents, we segment the procedure into three parts. We use one example to illustrate the procedure. The example is from a real-world Rug Pull incident called ZHONGHUA (Certik, 2024). Figure 1 shows a simplified code snippet of the scam token for illustration.
Figure 1. Code Snippet of ZHONGHUA Token
Phase1: Tactic Injection. Rug Pulls typically begin with scammers embedding fraudulent tactics to facilitate fraud ($S_{1}$: Fraudulent Tactics Injection), followed by simulating the launch of a legitimate token, termed as $S_{2}$: Token Launch. In $S_{1}$, scammers typically introduce hidden fraudulent tactics into the token. For instance, the token contract may contain malicious code that enables scammers to manipulate the token. Then, in $S_{2}$, scammers simulate the launch of a legitimate token. This step includes creating and promoting the new token to attract investors and establishing a liquidity pool by pairing the token with reputable assets to boost credibility.
Example: The scammer designated a tax-exempt address (line6 in Figure 1) controlled by the scammer within the token contract. Then they launched the token and locked the initial liquidity to deveive investors, promoting the token as a valuable investment.
Insights. A typical Rug Pull involves injecting malicious code into the token contract($S_{1}$) and launching the new token($S_{2}$), followed by inflating the token price to lure investors($S_{3}$), and extracting liquidity to drain valuable assets($S_{4}$). Phase1 involves embedding malicious code that facilitates control over the token and Phase2 includes suspicious transaction behaviors that simulate organic growth. Effective Rug Pull detection before reaching Phase3 relies on integrating code risk information from Phase1 and suspicious transaction behavior from Phase2, which are complementary. In real-world scenarios, Rug Pulls do not always exhibit both stages explicitly. Relying on only one aspect can result in potential false positives and false negatives. Therefore, an approach that integrates code and transaction information is essential for comprehensive Rug Pull detection.
3.3. Rug Pull Risk Feature
Table 1. Definition of Rug Pull Code Risks
@@ 3.3. Rug Pull Risk Feature
3.3.1. Rug Pull Code Risk
We analyze the source code within Rug Pull tokens, categorizing code risks into three major categories and eight subcategories. Specifically, we have identified Sale Restrict behaviors in 203 Rug Pull token contracts, Variable Manipulation in 143 contracts, and Balance Tamper in 189 contracts. Our analysis reveals the prevalence of these code risks in different Rug Pull schemes. Table 1 shows the definition of Rug Pull code risks.

### v2

# RPHunter: Unveiling Rug Pull Schemes in Crypto Token via Code-and-Transaction Fusion Analysis

@@ Abstract
In this paper, we propose RPHunter, a novel technique that integrates code and transaction for Rug Pull detection. First, RPHunter establishes declarative rules and performs flow analysis to extract code risk information, further constructing a semantic risk code graph (SRCG). Meanwhile, to leverage transaction information, RPHunter formulates dynamic token transaction activities as a token flow behavior graph (TFBG) in which nodes and edges are characterized from network structure and market manipulation perspectives. Finally, RPHunter employs graph neural networks to extract complementary features from SRCG and TFBG, integrating them through an attention fusion model to enhance the detection of Rug Pull. We manually analyzed 645 Rug Pull incidents from code and transaction aspects and constructed a ground-truth dataset. We evaluated RPHunter on our dataset, achieving a precision of 95.3%, a recall of 93.8% and an F1 score of 94.5%, which highlights superior performance compared to existing methods. Furthermore, when applied to the real-world scenarios, RPHunter has identified 4801 Rug Pull tokens, achieving a precision of 90.7%.
Index Terms:
Smart Contracts, Rug Pull, Static Analysis, Graph Neural Networks.
I Introduction
@@
Blockchain technology has revolutionized the financial landscape, enabling direct peer-to-peer transactions and financial services without centralized intermediaries. A diverse range of decentralized applications, including decentralized exchanges, lending platforms, and prediction markets, have been built on it. At the time of writing, the global crypto market cap has reached \$3.25 trillion [1]. However, the rapid expansion and the inherent openness of DeFi also pose considerable security challenges. Among these, Rug Pulls have emerged as a prominent form of scam, leading to significant financial losses, with approximately \$85.4 million in 2024 alone [2].
@@
Rug Pulls are a pernicious form of scam prevalent within the DeFi ecosystem, where scammers lure investors with seemingly promising tokens. Once a substantial amount of funds is collected, these scammers abruptly withdraw the funds and abandon the project entirely, leaving investors with worthless tokens. To detect such scams, researchers have proposed various methods [3, 4, 5, 6, 7], broadly classified into rule-based methods [4, 5] and learning-based methods [6, 7]. Rule-based methods typically use static analysis to identify vulnerabilities in token contracts, requiring domain knowledge of known Rug Pull patterns. Despite their promise in detecting these known patterns, these methods are prone to high false negatives due to their limited detection scope. In contrast, learning-based methods employ statistical features from token transactions and opcode data to train machine learning models for detection. However, these methods struggle to identify coordinated manipulation involving multiple addresses and detect Rug Pulls before they occur, suffering from high false positives and poor practicality.
@@
The continuous evolution of real-world Rug Pulls makes them increasingly complex and diverse, posing significant challenges to existing methods. (1) Real-world Rug Pull schemes are constantly evolving and it is a complex interplay between malicious code and suspicious behaviors, yet there is still a lack of research that investigates the underlying mechanisms. Although some studies [4, 5] have partially uncovered the mechanisms, a comprehensive empirical study focusing on the mechanisms of real-world Rug Pull cases is still missing. (2) In real-world scenarios, Rug Pulls are characterized by a complex interplay between malicious code logic and suspicious transaction behaviors. However, these two threat vectors do not always co-occur in practice. Some schemes exploit only covert code-level traps, while others rely primarily on transaction-level manipulation such as wash trading. Relying solely on code or transaction information to detect Rug Pulls can lead to false negatives and false positives. However, existing methods fail to organize and leverage information from both code and transaction perspectives.
In this paper, we focus specifically on Rug Pull schemes occurring on EVM-compatible blockchains. To address the aforementioned challenges, we conducted an empirical study to uncover the underlying mechanisms of Rug Pull. Building upon prior studies [5, 4], which analyzed 103 and 201 samples respectively, we construct a comprehensive real-world dataset on EVM blockchains from security platforms, comprising 645 real-world Rug Pull incidents. Each case is manually examined from both code and transaction dimensions, enabling us to summarize the common Rug Pull procedure and identify key risk features.
@@
We evaluated RPHunter on our ground-truth dataset. We compared the performance of RPHunter with five state-of-the-art methods. The results demonstrate that RPHunter detects Rug Pull tokens with a precision of 95.3%, a recall of 93.8%, and an F1 score of 94.5%, significantly outperforming existing methods. Additionally, to assess its real-world applicability, we applied RPHunter to newly created Ethereum Mainnet crypto tokens. RPHunter identified 4801 Rug Pull tokens. We randomly sample tokens with a 99% confidence level and an 8% confidence interval for manual verification. The results show that RPHunter achieves an overall precision of 90.7%. A prototype of RPHunter, all of our datasets, and experimental results are available online [8].
@@
  • We construct a comprehensive manually analyzed Rug Pull dataset, comprising 645 Rug Pull incidents, and summarize the typical Rug Pull procedure along with risk features from code and transaction perspectives.
@@
  • We propose RPHunter, a novel method that integrates the complementary information from malicious code risk and suspicious transaction behavior, utilizing GNNs for graph feature learning to detect Rug Pull schemes.
@@
  • Experimental results show that RPHunter outperforms the current SOTA methods on our dataset with a precision of 95.3%, a recall of 93.8%, and an F1 score of 94.5%. Additionally, we have identified 4801 Rug Pull tokens in real-world crypto tokens, achieving a precision of 90.7%.
II Background
II-A Smart Contract and Transaction
@@
In this paper, we focus on blockchain networks running on the Ethereum Virtual Machine (EVM), which adopt an account-based model. Smart contracts are self-executing programs with predefined rules and logic encoded in their code, which are typically written in a Turing-complete programming language (e.g., Solidity, Viper). To deploy smart contracts on the blockchain, they need to be compiled into bytecode and then submitted to the blockchain with transactions. Once deployed, smart contracts are immutable and can be accessed through public functions, serving as entry points for interactions with external owned accounts (EOAs) or other contracts. Transactions are essential as the medium for facilitating interactions within the blockchain network, acting as the vehicle for operations such as contract creation and invocation.
II-B Token and Token Standard
@@
Tokens are a kind of cryptocurrency implemented by smart contracts which manage internal book-keeping to record the information of token holders (e.g., their addresses and shares), and support token-related activities such as token balance inquiry and token transfer. Token standards specify a unified protocol for token creation, distribution, and interaction, ensuring compatibility and functionality across various platforms. A token smart contract should implement standard interfaces and emit standard events defined by token standards, otherwise, the token cannot be purchased or sold on the blockchain. For example, the ERC-20 standard defines 6 standard interfaces and 2 standard events for transferring tokens, retrieving account balances, and approving the use of tokens, alongside events for logging transactions.
II-C Rug Pull
@@
Rug Pulls are a prevalent form of scam within the DeFi ecosystem, frequently executed on Decentralized Exchanges (DEXs). Unlike centralized exchanges, which match buyers and sellers via orderbook, DEXs facilitate trades through liquidity pools. Liquidity pools are a fundamental mechanism in DEXs where tokens are paired with reputable assets (e.g., WETH), allowing users to trade tokens directly from the pool without intermediaries. The liquidity pool dynamically adjusts token prices based on the principles of supply and demand.
In a legitimate token launch, the process typically begins with the deployment of a token contract, followed by the creation of a liquidity pool pairing the new token with a reputable asset. The token issuer then receives liquidity provider (LP) tokens representing their share of the pool. To build trust and long-term value, responsible developers often lock these LP tokens in smart contracts (e.g., via time-locks) to prevent sudden withdrawal of liquidity. To generate revenue, developers may promote their token by associating it with unique features or utility (e.g., staking rewards), and profit through mechanisms such as transaction fees or other incentive models. In contrast, a Rug Pull is a malicious scheme where mimic the initial steps of a token launch—deploying a token contract and creating a liquidity pool—but with hidden intentions. They may embed malicious code in the contract or simply choose not to lock the initial liquidity. As the token’s price rises, the scammers exploit their unrestricted control over the liquidity pool or token contract. They swiftly exchange a large quantity of scam tokens for the paired reputable asset, effectively draining the pool. This sudden removal of liquidity leaves remaining investors holding worthless tokens with no market for resale.
III Real-world Rug Pull Study
III-A Data Collection and Analysis
III-A1 Rug Pull Incidents Collection
@@
Existing studies [5, 4] have provided preliminary insights into Rug Pull schemes, analyzing 103 and 201 Rug Pull samples, respectively. To enable a more comprehensive empirical study, we significantly expanded upon these datasets by collecting Rug Pull incidents from blockchain security platforms including De.Fi [9], Peckshield [10], Certik [11], SlowMist [12] and Beosin [13]. Ultimately, we collected a total of 1048 Rug Pull incidents that occurred before May 2024, encompassing multiple blockchain networks, including ETH [14], BSC [15], FTM [16], BASE [17], Arbitrum [18] and Polygon [19].
Figure 1: ZHONGHUA Rug Pull Procedure
@@
After analyzing these Rug Pull incidents, we excluded 131 tokens without source code, hindering our precise analysis of these cases, and 272 tokens without transaction history, which suggests that these projects did not result in actual Rug Pulls. Finally, we obtained 645 Rug Pull incidents and constructed a comprehensive manually analyzed Rug Pulls dataset [8]. While some studies [6, 7] have constructed datasets of thousands of Rug Pull cases, they have two limitations: (1) these cases lack manual analysis and verification (e.g., code risk analysis), (2) rely on heuristic rules to select these cases, making them susceptible to false positives.
III-A2 Incidents Analysis
@@
To better understand Rug Pull schemes, we conducted a comprehensive manual analysis of the incidents. For each Rug Pull incident, we conducted a detailed examination of the tokens, focusing on both the source code and the transaction history to uncover the underlying mechanisms of each Rug Pull. To ensure a high level of accuracy, we recruited three experts with two years of experience in smart contract security research.
Specifically, we adopted an open card sorting [20] approach to categorize and analyze these incidents. For each Rug Pull case, we created a card comprising three components: (1) the incident name, (2) the description derived from the original incident report, and (3) the identified root cause. The root cause analysis involved two dimensions code and transaction. We performed the manual analysis, following the detailed steps illustrated in [21]. Each expert independently examined the source code of the token contract to identify and categorize suspicious code patterns (e.g., backdoor functions). Then using CoinMarketCap [1] historical data, experts reviewed each token’s price trajectory to identify the precise transaction that triggered a rapid collapse in value, then analyzed the preceding transaction behaviors. After completing their independent assessments, the experts documented the identified root causes. A consensus-based voting session was then held to resolve discrepancies through discussion and determine the most plausible root cause for each incident. We collaboratively reviewed the identified root causes, formulating a generalized Rug Pull procedure (Section III-B), which captures the essence of these frauds. Additionally, we further summarized the Rug Pull risk features that emerge from code risk (Section III-C1) and transaction behavior (Section III-C2) two perspectives.
TABLE I: Definition of Rug Pull Code Risks
@@
III-B Rug Pull Procedure
After a systematic analysis of the Rug Pull procedure based on these real-world incidents, we segment the procedure into three parts. We use one example to illustrate the procedure. The example is from a real-world Rug Pull incident called ZHONGHUA [22]. Fig. 1 shows the simplified procedure of the scam token for illustration.
Phase1: Tactic Injection. Rug Pulls typically begin with scammers embedding fraudulent tactics to facilitate fraud ($S_{1}$: Fraudulent Tactics Injection), followed by simulating the launch of a legitimate token, termed as $S_{2}$: Token Launch. In $S_{1}$, scammers typically introduce hidden fraudulent tactics into the token. For instance, the token contract may contain malicious code that enables scammers to manipulate the token. Then, in $S_{2}$, scammers simulate the launch of a legitimate token. This step includes creating and promoting the new token to attract investors and establishing a liquidity pool by pairing the token with reputable assets to boost credibility. To further deceive investors, they may lock the initial liquidity, leading investors to believe that it is secure and irreversible.
Example: The scammer designated a tax-exempt address (line6 in Fig. 1) controlled by the scammer within the token contract. Then they launched the token and locked the initial liquidity to deveive investors, promoting the token as a valuable investment.
Summary and Insights. While the above procedure outlines a generalized lifecycle of typical Rug Pull schemes-begining with the injection of malicious tactics ($S_{1}$) and the new token launch($S_{2}$), followed by inflating the token price to lure investors($S_{3}$), and extracting liquidity to drain valuable assets($S_{4}$), our empirical analysis reveals that not all Rug Pull strictly adhere to this full sequence. In particular, we observed two notable variations:
  • In some cases, scammers rely solely on malicious logic embedded in the token contract, such as backdoor functions that allow unrestricted minting. They do not engage in artificial price inflation, but instead wait for organic token appreciation before silently exploiting the backdoor.
  • Other cases show that scammers can perform Rug Pulls without introducing malicious code. Instead, they deliberately avoid locking all initial liquidity, allowing them to silently withdraw the unlocked liquidity at an opportune moment without invoking exploitative code.
These variations highlight the heterogeneity of Rug Pull schemes. A key insight from our study is that effective Rug Pull detection, especially before the scheme reaches the irreversible Phase 3, requires the integration of both code-level risk indicators from Phase 1 and suspicious transaction behaviors from Phase 2. Focusing exclusively on one dimension may lead to potential false positives and false negatives. Therefore, we aim to propose a hybrid approach on EVM-compatible blockchains by jointly analyzing token contract code and early-stage transactional behaviors, enabling early-stage Rug Pull detection across heterogeneous scam patterns.
III-C Rug Pull Risk Feature
III-C1 Rug Pull Code Risk
We analyze the source code within Rug Pull tokens, categorizing code risks into three major categories and eight subcategories. Specifically, we have identified Sale Restrict behaviors in 203 Rug Pull token contracts, Variable Manipulation in 143 contracts, and Balance Tamper in 189 contracts. Our analysis reveals the prevalence of these code risks in different Rug Pull schemes. Table I shows the definition of Rug Pull code risks.